Tiger Golf Accessories sells golf shoes, gloves, and a laser-guided range-finder that measures distance. Shown below are unit cost and sales data.

Pairs of
Shoes

Pairs of
Gloves

Range-
Finder

Unit sales price $106 $30 $245
Unit variable costs 61 11 202
Unit contribution margin $45 $19 $43
Sales mix 33 % 41 % 26 %


Fixed costs are $652,726.

Verify that the mix of sales units to be sold at the break-even point for each product line will generate a zero net income by creating a CVP INCOME STATEMENT

Tiger Golf Accessories
CVP Income Statament
$
$
$
$
Answer #1

Required:

  • Total weighted-average unit contribution margin = ($45 * 33%) + ($19 * 41%) + ($43 * 26%)

= $14.85 + $7.79 + $11.18

= $33.82

  • Break-even point = Fixed cost / weighted-average unit contribution margin

= $652,726 / $33.82

= 19,300 units.

  • Number of units sold at Break-even point

Shoes = (19,300 * 33%) = 6,369 units

Gloves = (19,300 * 41%) = 7,913 units

Range finder = (19,300 * 26%) = 5,018 units.

  • Income statement:
Sales of shoes (19,300 * 33%) * $106 $675,114
Sales of gloves (19,300 * 41%) * $30 $237,390
Sales of range finder (19,300 * 26%) * $245 $1,229,410
Total sales $2,141,914
(-) Variable cost:
(6,369*61) + (7,913*$11) + (5,018*$202) ($1,489,188)
Contribution margin $652,726
(-) Fixed cost ($652,726)
Net income $0